OpenStudy (anonymous):

x/2+2/3=5/6 how to get the value of X

OpenStudy (anonymous):

x/2 = 5/6 - 2/3 = 5/6-4/6 =1/6 ===>x=2/6=1/3

OpenStudy (anonymous):

In soliving Multi-step Equations

OpenStudy (anonymous):

get L.C.M X/2=5/6-2/3 X/2=(5-4)/6 X/2=1/6 X=1/6.2 X=1/3